OpenStudy (anonymous):

Divide the following polynomials, then place the answer in the proper location on the grid. Write the answer in descending powers of a. Divide 5a2 + 6a - 9 into 25a4 + 19a2 - a - 78.

OpenStudy (saifoo.khan):

into means multiply!

OpenStudy (saifoo.khan):

5a2 + 6a - 9 --------------- 25a4 + 19a2 - a - 78. this?

OpenStudy (anonymous):

Just tell us in which part u r having no problem, makes no sense if we do all ur problems

OpenStudy (saifoo.khan):

hellO? is it like this??? :S

OpenStudy (anonymous):

Is it divide by or divide into?

OpenStudy (saifoo.khan):

same confusion.

OpenStudy (anonymous):

\[\frac{25 a^4+19 a^2-a-78}{5 a^2+6 a-9}=5 a^2-6 a+20-\frac{175 a-102}{5 a^2+6 a-9} \]